OpenStudy (anonymous):

5x + 9y = -54 2x + y = 20 a.) Find the solution to the system of equations using either substitution of the linear combination method. Show all your work. I've tried but I can't seem to understand how to do this? Does anybody mind giving me steps on how to do either method?

OpenStudy (hockeychick23):

so lets eliminate y by multiplying the second equation by 9: 5x+9y= -54 (2x+y= 20)9= 18x+9y= 180 So now we have these two equations: 5x + 9y= -54 18x+ 9y= 180 Lets now subtract the second equation from the first: -13x= - 234 to get x alone lets divide both sides by -13 -13x/-13= -234/-13 x=18 Now we can plug x into either problem: 2(18)+y=20 36+y= 20 Now we subtract 36 from both sides. y= 20-36 y= -16